A man walks on a straight road from his home to a market 2.5 km away with a speed of 5 km/hr. Finding the market closed, he instantly turns and walk back home with a speed of 7.5km/her. The average speed of the man over the interval of time 0 to 40 mins is equal to____
1
Expert's answer
2021-03-14T19:19:52-0400

The average speed can be found as follows:


"v_{avg}=\dfrac{d_{tot}}{t_{tot}},""v_{avg}=\dfrac{d_1+d_2}{t_1+t_2}."

The distance traveled to the market:


"d_1=2.5\ km."

Let's find time taken to go to the market:


"t_1=\dfrac{d_1}{v_1}=\dfrac{2.5\ km}{5\ \dfrac{km}{h}}=\dfrac{1}{2}\ h=30\ min."

Let's find time taken to return back:


"t_2=\dfrac{d_2}{v_2}=\dfrac{2.5\ km}{7.5\ \dfrac{km}{h}}=\dfrac{1}{3}\ h=20\ min."


But, in fact, we need his average speed in the interval of time 0 to 40 minutes, so "t_2=10\ min".

Let's find the distance traveled in 10 minutes while the man returning back:


"d_2=v_2t_2=7.5\ \dfrac{km}{h}\cdot10\ min\cdot\dfrac{1\ h}{60\ min}=1.25\ km."

Finally, we can find the average speed of the man over the interval of time 0 to 40 minutes:


"v_{avg}=\dfrac{2.5\ km+1.25\ km}{40\ min\cdot\dfrac{1\ h}{60\ min}}=5.625\ \dfrac{km}{h}."